Circular No. 6167

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

MARKARIAN 421 T. Takahashi, F. Makino, and T. Kii, Institute of Space and Astro-

nautical Science; M. Tashiro, H. Kubo, and K. Makishima, University of Tokyo; T. Ohashi and N. Y. Yamasaki, Tokyo Metropolitan University; and G. Madejski, Goddard Space Flight Center, NASA, write: “The BL Lac object Markarian 421 was observed from the x-ray astronomy satellite ASCA during Apr. 25.8–26.1 UT. The preliminary energy flux obtained from GIS and SIS abroad ASCA was 1.8 × 10^{−10} erg cm^{−2} s^{−1} (2–10 keV), which is almost the same as that observed in the previous x-ray/TeV flare in May 1994 (IAUC 5993, 5996). The photon index was ∼ 2.5. Further x-ray observations were planned for Apr. 28, 29, May 1, 3, 5, 6, and 8. Observations at optical, infrared, radio, and other wavelengths are encouraged. For more details of the x-ray observation schedule, contact T. Takahashi

(e-mail takahasi@astro.isas.ac.jp).”

J. Buckley, J. Quinn, and T. C. Weekes, Smithsonian Astrophysical Observatory; M. Catanese, D. A. Carter-Lewis, F. Krennrich, R. C. Lamb, and J. Zweerink, Iowa State University; C. W. Akerlof, D. Meyer, and M. Schubnell, University of Michigan; J. Gaidos, G. Sembroski, and C. Wilson, Purdue University; D. J. Fegan, R. Lessard, and J. McEnery, University College, Dublin; M. F. Cawley, St. Patrick’s College, Maynooth; and S. Biller, A. M. Hillas, A. Rodgers, and J. Rose, University of Leeds, communicate: “Markarian 421 has been in a period of unusually high activity in TeV γ-rays. An intense observing campaign with the Whipple Gamma Ray Telescope was timed to bracket the EGRET observations that began on Apr. 25. Analysis of nightly observations gives the following measurements: Apr. 21, 1.12 ± 0.13 photons/min; 22, 1.18 ± 0.11; 23, 0.84 ± 0.08; 24, < 0.44; 25, 0.42 ± 0.11; 26, 1.90 ± 0.12; 27, 1.36 ± 0.11; 28, 0.81 ± 0.12; 29, 1.15 ± 0.10; 30, 1.05 ± 0.14; May 1, 0.68 ± 0.12; 2, < 0.76. Observations generally began at 3^{h}30^{m} UT each day and continued for 3 hr. The statistical significance of the detections ranged from 3.6σ (Apr. 25) to 15.4σ (Apr. 26); upper limits are quoted at the 3σ level. The average photon flux between 1995 Jan. 1 and Apr. 5 was also variable but averaged 0.9 photon/min, corresponding to a flux of 4.3 × 10^{−11} photon cm^{−2} s^{−1}. The energy threshold of all observations was set at 300 GeV in this preliminary analysis. Observations at other wavelengths are encouraged.”

1995 May 4 Daniel W. E. Green